# {strip}[Uncaptioned image]

### v1

@@ Abstract.
VOS is a first-of-its-kind instructional OS that: (1) Runs on commodity, portable hardware. (2) Showcases modern features, including per-app address spaces, threading, commodity filesystems, USB, DMA, multicore, self-hosted debugging, and a window manager. (3) Supports rich applications such as 2D/3D games, music and video players, and a blockchain miner.
Unlike traditional instructional systems, VOS emphasizes strong motivation for building systems—supporting engaging, media-rich apps that go beyond basic terminal programs. To achieve this, we design VOS to strike a careful balance between essential OS complexity and overall simplicity. Our method, which we call inverse engineering, breaks down a full-featured OS into a set of incremental, self-contained prototypes. Each prototype introduces a minimal set of OS mechanisms, driven by the needs of specific apps. The construction process (i.e., forward engineering) then progressively enables these apps by bringing up one mechanism at a time.
VOS makes it accessible for a wider audience to experience building a software system that is self-contained and usable in everyday scenarios.
@@ 1. Introduction
Beyond these technical concepts, OS courses are also expected to offer a unique experience: building a working system that students can both demonstrate and use themselves. To this end, the OS course is the apex of an undergraduate CS curriculum – where students integrate their knowledge to construct a complete computer system.
1.1. A Missed Opportunity
Unfortunately, we observe a declining interest in OS—and systems software in general—among college students. Several factors contribute to this trend: the dominance of AI, the perception that systems work is overly complex, and the frustration often associated with modern systems software. One telling indicator is that many premier North American universities no longer require an OS course for undergraduate CS majors (see Section 2).
@@ 1.2. Principles
Our response is an instructional OS called Proto, which embodies a guided process for modern OS building.
Our first priority is to establish clear and compelling motivations for system building — motivations that go beyond running simple terminal programs. To this end, we target (P1) appealing apps: engaging, interactive programs that ideally evoke an emotional response from the OS builders themselves. This motivation directly influences our design choices, such as treating the framebuffer as a first-class peripheral and prioritizing support for USB and audio IO. We further emphasize (P2) demonstrability: The OS should be able to run on real commodity hardware in addition to emulators. Ideally, it should be portable enough to carry around for live demonstrations, enabling students to showcase their work long after the course ends.
Our second consideration is managing complexity: balancing the technical depth needed to support engaging apps with the limited effort students can realistically invest (typically 17 weeks and no more than 150 hours). To achieve this, we adopt (P3) incremental prototyping: The OS is delivered to students in multiple self-contained snapshots, each augmenting earlier ones and is fully functional. Compared to a “big bang” approach, incremental prototypes provide clear baselines for continued development. More broadly, incremental prototyping is a proven method in complex engineering projects such as the Ford Model T, Boeing 747, and SpaceX rockets (burchardt2020development, ; wu2012rapid, ; spaceX, ). We also require (P4) minimum viable implementations: Every new OS feature—whether it’s a syscall, threading, or multicore scheduling—must directly support a target app or an essential component of it. We will evaluate the efficacy of these principles in Section 6.
@@ 1.3. Design
Through a process dubbed inverse engineering, we first fully implement VOS, before decomposing it into five incremental snapshots, referred to as “prototypes 1–5”. Further through inverse engineering, we map each prototype to one or multiple target applications, ranging from simple 3D animations (Prototype 1) to video games (Prototype 3) and video playback (Prototype 5). See Table 1 for prototype features and their target apps.
The forward engineering (i.e. OS construction by students) is then driven by enabling one app after another, re-architecting earlier prototypes and adding new OS features including per-app address spaces, multimedia output, USB peripherals, filesystems, DMA, and multicore execution. At a higher level, this process resembles interactive designs — akin to video games or theme park experiences — by keeping students as OS builders motivated with clear, meaningful goals, rewarding them with tangible progress (making apps functional), and maintaining engagement through timely transitions between different prototype stages.
@@ 1.4. Results & Contributions
First, VOS is lightweight. Its kernel core is less than 10K SLoC, comparable to existing instructional OSes (§6.1).
Second, VOS is demonstration-ready. It runs on both software emulators and low-cost ARMv8 hardware. As shown in Figure 1, VOS supports a variety of apps: it can run games (e, i), play music and videos (g, j), and even project presentation slides for classroom use (f).
Third, VOS is fast and efficient. Through a careful balance of optimization and simplicity, VOS achieves strong performance on low-cost ARMv8 platforms. On kernel microbenchmarks (§6.2), it incurs overheads similar to xv6 (cox2011xv6, ); it performs comparably to Linux and FreeBSD on most benchmarks. On app benchmarks (§6.3), VOS runs DOOM at around 60 FPS and plays back video at around 26 FPS – on par with Linux and FreeBSD running the same apps on the same hardware. VOS also scales to quad-core processors. When running on a handheld device (Figure 1(h)), the system typically draws less than 4 watts and operates for several hours on a single battery charge (§6.4).
Finally, VOS is pedagogically effective. Our user study shows that most students found the design principles (P1–P4) directly contributed to their learning experience and sparked enthusiasm for systems software (§6.5).
We made the following contributions:
  • Methodologies. This paper presents a case of inverse engineering: first designing a modern OS, then systematically decomposing it into a series of incremental, self-contained prototypes, and mapping each subset of OS mechanisms to target apps that critically depend on them.
  • Design Choices. We report the architectures chosen for each prototype, balancing trade-offs among app features, OS complexity, and modern OS concepts to be conveyed.
  • Software Artifacts. Developed from scratch, VOS is the first-of-its-kind instructional OS that: (1) Runs on commodity, portable hardware. (2) Showcases modern features, including per-app address spaces, threading, commodity filesystems, USB, DMA, multicore, self-hosted debugging, and a window manager. (3) Supports rich applications such as 2D/3D games, music and video players, and a blockchain miner.
The code is publicly available at: {anonymized for review}.
@@ 2. Background
2.1. Context
The authors’ institution recently made OS elective for CS majors. Similarly, among the top 30 North American universities (csranking, ), most (20) no longer require OS as a core course. This shift presents both a wake-up call and an opportunity for OS educators: (1) Instructional OS must stay relevant by appealing to broader students, highlighting its practical and demonstrable value. (2) As an elective following Computer System Organization, the OS course now addresses more motivated, better-prepared students. AI tools (e.g., CoPilot (msft_copilot, )) further empower them to take on complex code.
@@ 2.2. Relation to existing OSes
Most instructional OSes are ”headless,” i.e., they lack GUI and instead support shell-based apps and utilities backed by simple file systems. Notable academic projects include Xinu (comer2015operating, ), xv6 (cox2011xv6, ), Nachos (christopher1993nachos, ), GeekOS (hovemeyer2004running, ), Pintos (pfaff2009pintos, ), MiniOS (otero2018minios, ), and the more recent EGOS-2000 (githubGitHubYhzhang0128egos2000, ). The open-source community also contributes instructional OSes (osfs00, ; raspberry-pi-os, ). Among these, only a small subset can run on actual hardware, typically on ARMv7 (francis2018raspberry, ; githubGitHubLdBECMXinu, ), ARMv8 (raspberry-pi-os, ; githubGitHubHongqinLirpios, ), and RISC-V (gebhard2024using, ; cox2011xv6, ; githubGitHubYhzhang0128egos2000, ).
While prior works—especially xv6—have greatly influenced our system, their designs heavily focus on conveying OS concepts, overlooking appealing apps (P1) and demonstrability (P2), which are central to our approach.
@@
A recent survey (ebling2024resources, ) suggests that more than half of participating universities use Linux for teaching OS. Other choices of production OSes include Android (andrus2012teaching, ), RT-thread (rt-thread, ), and RISCOS (riscos, ). Students typically modify parts of these systems, such as CPU schedulers. This approach falls short for several reasons: (1) It does not align with our goal of providing an end-to-end system-building experience. Modifying a large, existing OS does not offer the same educational value as building one from scratch. (2) Production OSes are not suitable for stripping down due to their inherent complexity and generality. They support a wide range of apps and hardware, resulting in excessive layers of indirection. (3) The relevance of Linux/Android kernel development may be limited for a broader audience.
@@
We are inspired by recent educational initiatives that guide students in constructing end-to-end systems capable of running graphical apps (one_student_one_chip, ; project-n, ). Compared to them: (1) We target commodity hardware as the primary platform. This choice significantly influences how we balance practical realism and simplicity, handle hardware quirks, and support debugging (see Section 3). (2) We emphasize demonstration-ready, usable apps, which in turn motivate support for modern OS features such as commodity filesystems, USB, and DMA. (3) These goals further necessitate our approach of inverse engineering.
2.3. Hardware Considertations
@@
We choose the Raspberry Pi 3B (Pi3) as our sole platform for several reasons: (1) Popular ISA: ARMv8 is simpler than x86, yet widely used (IoT, phones, Macs), with extensive documentation and relevance; (2) Availability: Released in 2016, Pi3 has sold over 23 million units, is priced at \$35, and will remain in production through 2029 (tomshardware2024raspberry, ); (3) Documentation: Pi3 is well-documented, with strong open-source support and an active community (raspberrypiRaspberryRaspberry, ); (4) Performance: Four Cortex-A53 cores at 1 GHz, NEON SIMD, and 1 GB of memory make Pi3 suitable for a range of gaming and media apps; (5) Peripheral Ecosystem: Pi3 is compatible with many expansion boards (called HATs (githubGitHubRaspberrypihats, )), including a Game HAT (see Section 6), which adds a 640p screen, battery, buttons, and speakers.
@@
3. VOS Overview
We next describe VOS in its complete form, which includes all the OS capabilities. This implementation is the starting point for our decomposition into a series of prototypes.
@@ 3. VOS Overview
  • donut. one or more spinning 3D tori (donut, ). comprise either textual characters (for UART text output) or pixels (for framebuffer).
  • mario, the LiteNES emulator (liteNes, ) supporting limited games, including Mario (1983).
  • sysmon. A floating, transparent window that visualize real-time CPU and memory usage.
  • slider. Player of slides in BMP, PNG, and GIF format. Intended for the OS builders to present their design.
  • DOOM. A famous 3D game that has been ported to virtually ”anything with a screen” (itrunsdoom, ). We build on the ”doom generic” (doom_generic, ) project and implement key event polling and texture/renderer management. We chose not to implement sound mixing as it would require a sound mixer.
  • music player. Intended to enhance user engagement and evoke emotions. It plays OGG files while displaying album covers. The importance of sound in games/computer programs and for tangible user experience is crucial (guillen2021role, ; kenwright2020there, ; robb2017impact, ).
  • video player. Playback of MPEG-1 videos.
  • launcher. A GUI frontend for launching programs; displaying an animated background.
  • blockchain. A multithreaded program for mining blocks.
Of these apps, a number of them were derived from the project-N(project-n, ). We also ported all console apps from xv6, including shell (enhanced with script execution) and utilities as ls, cat, and echo. Most user apps are in C and only a few are in C++ (blockchain).
@@
VOS is a monolithic kernel with an architecture similar to xv6. The kernel and user apps run at EL1 and EL0 of ARMv8, respectively. The OS supports user processes and both user and kernel threads, and can execute them across all four CPU cores of the Pi3. The kernel is written in a mix of C and ARMv8 assembly. It includes CPU-specific code for the ARM Cortex-A53 and board-specific device drivers for the Pi3.
VOS uses virtual memory with separate address spaces for each user app. User space starts at 0x0; kernel space uses addresses prefixed with 0xffff. The kernel maps physical memory and I/O in 1 MB blocks, while user apps use 4 KB pages. Only the user stack uses demand paging.
VOS implements 28 syscalls across three categories: task management (e.g., fork, sleep), file system (e.g., open, close), and threading/synchronization (e.g., clone, semaphore). The kernel exports device files (e.g., /dev/fb for framebuffer and /dev/events for keyboard) and proc files (e.g., /proc/cpuinfo and /proc/meminfo).
@@
VOS supports a selective set of IO peripherals, tailored for the target apps: framebuffer (for HDMI output), USB keyboards, GPIO (for Game HAT buttons), PWM (for sound output via a 3.5mm jack), SD cards, and UART. The decision to include USB is a careful balance between demonstrability and code simplicity, as discussed in Section 4.4.
VOS implements a window manager (see a photo in Figure 1(m)) allowing multiple apps to render to the screen and for input events to be dispatched to the app with focus.
@@
On the Pi3, VOS runs from an SD card with two partitions. Partition 1 contains the kernel image, which is loaded and launched by the Pi3 firmware. The kernel image packs the kernel code/data as well as an opaque ramdisk dump. The ramdisk, which is formatted with the ext2-like filesystem, incorporates all the user programs as ELF executables. Partition 2, formatted as FAT32, stores user files such as videos and game assets. At boot time, the kernel mounts its primary filesystem on the ramdisk and the secondary filesystem on the SD card’s partition 2. This setup allows users to conveniently exchange files between VOS and their personal devices (computers, phones, cameras) by copying files to and from partition 2. See Section 4.5 for details and rationale for supporting FAT32.

### v2

# Proto: A Guided Journey through Modern OS Construction

@@ Abstract.
Proto is a new instructional OS that runs on commodity, portable hardware. It showcases modern features, including per-app address spaces, threading, commodity filesystems, USB, DMA, multicore support, self-hosted debugging, and a window manager. It supports rich applications such as 2D/3D games, music and video players, and a blockchain miner. Unlike traditional instructional systems, Proto emphasizes engaging, media-rich apps that go beyond basic terminal programs. Our method breaks down a full-featured OS into a set of incremental, self-contained prototypes. Each prototype introduces a minimal set of OS mechanisms, driven by the needs of specific apps. The construction process then progressively enables these apps by bringing up one mechanism at a time.
Proto enables a wider audience to experience building a self-contained software system used in daily life.
@@
journalyear: 2025copyright: ccconference: ACM SIGOPS 31st Symposium on Operating Systems Principles; October 13–16, 2025; Seoul, Republic of Koreabooktitle: ACM SIGOPS 31st Symposium on Operating Systems Principles (SOSP ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3731569.3764811isbn: 979-8-4007-1870-0/25/10ccs: Software and its engineering Operating systems
@@ 1. Introduction
Beyond these technical concepts, OS courses are also expected to offer a unique experience: building a working system that students can both demonstrate and use themselves. To this end, the OS course is the apex of an undergraduate CS curriculum, where students integrate their knowledge to construct a complete computer system.
Unfortunately, we observe a declining interest in OS—and systems software in general—among college students that we teach, owing to several factors: the dominance of AI (DominanceOfAI, ), the perception that systems work is overly complex, and the frustration often associated with modern systems software. One telling indicator is that many premier North American universities  (csranking, ) (20 out of top 30) no longer require OS as a core course for undergraduate CS majors. This shift presents both a wake-up call and an opportunity for OS educators: (1) Instructional OS must stay relevant by appealing to broader students, highlighting its practical and demonstrable value. (2) As an elective following Computer System Organization, the OS course now addresses more motivated, better-prepared students. AI tools (e.g., Copilot (msft_copilot, )) further empower them to take on complex code.
@@ 1.1. Principles
Our response is an instructional OS called Proto, which manifests a guided journey through building an OS with rich applications.
Our first priority is to establish clear and compelling motivations for system building—motivations that go beyond running simple terminal programs. To this end, we first target (P1) appealing apps: building engaging, interactive programs that ideally evoke an emotional response from the OS builders themselves. This motivation directly influences our design choices, such as treating the framebuffer as a first-class peripheral and prioritizing support for USB and audio IO. We further emphasize (P2) demonstrability. The OS should be able to run on real commodity hardware in addition to emulators. Ideally, it should be portable enough to carry around for live demonstrations, enabling students to showcase their work long after the course ends.
Our second consideration is managing complexity: balancing the technical depth needed to support engaging apps with the limited effort students can realistically invest (typically 17 weeks and no more than 150 hours). To achieve this, we adopt (P3) incremental prototyping. The OS is delivered to students in multiple self-contained snapshots, each augmenting earlier ones and is fully functional. Compared to a “big bang” approach, incremental prototypes provide clear baselines for continued development. We note that a number of prior instructional OSes were built with incremental prototyping (JOS, ; chickadee, ; pfaff2009pintos, ). More broadly, incremental prototyping is a proven method in complex engineering projects such as the Ford Model T, Boeing 747, and SpaceX rockets (burchardt2020development, ; wu2012rapid, ; spaceX, ). It is also common to many instructional OSes. We also require (P4) minimum viable implementations. Every new OS feature—whether it’s a syscall, threading, or multicore scheduling—must directly support a target app or an essential component of it. We will evaluate the efficacy of these principles in Section 7.
@@ 1.2. Design
We first fully implement Proto, before decomposing it into five incremental snapshots, referred to as “prototypes 1–5”. Further, we map each prototype to one or multiple target applications, ranging from simple 3D animations (Prototype 1) to video games (Prototype 3) and video playback (Prototype 5). See Table 1 for prototype features and their target apps.
The OS construction by students is then driven by enabling one app after another, augmenting earlier prototypes (so that they build upon familiar designs) and adding new OS features including per-app address spaces, multimedia output, USB peripherals, filesystems, DMA, and multicore execution. At a higher level, this process resembles interactive designs—akin to video games or theme park experiences—by keeping students as OS builders motivated with clear, meaningful goals, rewarding them with tangible progress (making apps functional), and maintaining engagement through timely transitions between different prototype stages in an incremental fashion.
@@ 1.3. Results & Contributions
Through its design, Proto achieved the following goals.
First, Proto is lightweight. Its kernel core is less than 10K SLoC, comparable to existing instructional OSes (§7.1).
Second, Proto is demonstration-ready. It runs on both software emulators and low-cost ARMv8 hardware. As shown in Figure LABEL:fig:collage, Proto supports a variety of apps: it can run games (e, i), play music and videos (g, j), and even project presentation slides for classroom use (f).
Third, Proto is fast and efficient. Through a careful balance of optimization and simplicity, Proto achieves strong performance on low-cost ARMv8 platforms. On kernel microbenchmarks (§7.2), it incurs overheads similar to xv6 (cox2011xv6, ); it performs comparably to Linux and FreeBSD on most benchmarks. On app benchmarks (§7.3), Proto runs DOOM at around 60 FPS and plays back video at around 26 FPS—on par with Linux and FreeBSD running the same apps on the same hardware. Proto also scales to quad-core processors. When running on a handheld device (Figure LABEL:fig:collage(h)), the system typically draws less than 4 watts and operates for several hours on a single battery charge (§7.4).
Finally, Proto is pedagogically effective. Our user study shows that most students found the design principles (P1–P4) directly contributed to their learning experience and sparked enthusiasm for systems software (§7.5).
In summary, this paper presents a series of incremental, self-contained OS prototypes motivated by appealing applications such as animations, media playback, and video games. Each prototype introduces a set of OS mechanisms explicitly mapped to target applications that critically depend on them. We report the architectures chosen for each prototype, balancing trade-offs among app features, OS complexity, and modern OS concepts to be conveyed. Our result is Proto, a new instructional OS that runs on commodity, portable hardware. It showcases modern features, including per-app address spaces, threading, commodity filesystems, USB, DMA, multicore, self-hosted debugging, and a window manager. It supports rich applications such as 2D/3D games, music and video players, and a blockchain miner. The code is publicly available at: https://github.com/fxlin/uva-os-main.
@@ 2. Background
This section positions Proto within the broader context of instructional OSes and explains our rationale for hardware choices.
@@ 2.1. Relation to existing OSes
Most instructional OSes are “headless”, i.e., they lack GUI and instead support shell-based apps and utilities backed by simple file systems. Notable academic projects include Embedded Xinu (gebhard2024using, ) (a port of Xinu (comer2015operating, )), xv6 (cox2011xv6, ), Nachos (christopher1993nachos, ), GeekOS (hovemeyer2004running, ), Pintos (pfaff2009pintos, ), MiniOS (otero2018minios, ), and the more recent EGOS-2000 (githubGitHubYhzhang0128egos2000, ). The open-source community also contributes instructional OSes (osfs00, ; raspberry-pi-os, ). Among these, only a small subset can run on actual hardware, typically on ARMv7 (francis2018raspberry, ; githubGitHubLdBECMXinu, ), ARMv8 (raspberry-pi-os, ; githubGitHubHongqinLirpios, ), and RISC-V (gebhard2024using, ; cox2011xv6, ; githubGitHubYhzhang0128egos2000, ).
While prior work—especially xv6—have greatly influenced our system, their designs heavily focus on conveying OS concepts, overlooking appealing apps (P1) and demonstrability (P2), which are central to our approach.
@@
A recent survey (ebling2024resources, ) suggests that more than half of participating universities use Linux for teaching OS. Other choices of production OSes include Android (andrus2012teaching, ), RT-thread (rt-thread, ), and RISCOS (riscos, ). Students typically modify parts of these systems, such as CPU schedulers. This approach falls short for several reasons. (1) It does not align with our goal of providing an end-to-end system-building experience. Modifying a large, existing OS does not offer the same educational value as building one from scratch. (2) Production OSes are not suitable for stripping down due to their inherent complexity and generality. They support a wide range of apps and hardware, resulting in excessive layers of indirection. (3) While working with production OSes like Linux/Android may excite some, their complexity may limit accessibility for a wider audience.
@@
We are inspired by recent educational initiatives that guide students in constructing end-to-end systems capable of running graphical apps (one_student_one_chip, ; project-n, ). Compared to them: (1) We target commodity hardware as the primary platform. This choice significantly influences how we balance practical realism and simplicity, handle hardware quirks, and support debugging (see Section 3). (2) We emphasize demonstration-ready, usable apps, which in turn motivate support for modern OS features such as commodity filesystems, USB, and DMA.
2.2. Student assignments
Next, we discuss student assignments built on prior educational OSes, including xv6 (cox2011xv6, ), Pintos (pfaff2009pintos, ), Embedded Xinu (embedded_xinu_project, ), GeekOS (hovemeyer2004running, ), EGOS-2000 (githubGitHubYhzhang0128egos2000, ), Chickadee  (chickadee, ), JOS (JOS, ), and WeensyOS (CS_61, ). Throughout, we use “lab” to refer to a single standalone submission from students.
Lab structure and granularity
Assignment design is largely shaped by the trade-off between learning objectives (breadth vs. depth) and the expected time commitment. Broadly, prior assignments fall into two types: breadth-oriented and depth-oriented. A breadth-oriented lab typically involves adding features to individual components (CS_61, ; hovemeyer2004running, ; black2009build, ; holland2002new, ), whereas a depth-oriented lab often requires implementing most or all of an OS subsystem (pfaff2009pintos, ; JOS, ; chickadee, ). Accordingly, breadth-oriented labs are usually independent (e.g., students receive fresh starter code for each lab), while depth-oriented labs are typically sequential, with each lab building on the codebase from previous labs (JOS, ; chickadee, ; pfaff2009pintos, ), culminating in a functional OS by semester’s end. Proto adopts a sequential approach in that labs build on each other, but is breadth-oriented in that each lab cuts across multiple OS components.
Lab topics
Although labs vary in granularity, each typically centers on a single OS topic—either a user component or a kernel subsystem. Common topics include: (1) shell implementation (JOS, ; CS_61_pset5, ); (2) CPU scheduling (otero2018minios, ; JOS, ; embedded_xinu_project, ; githubGitHubYhzhang0128egos2000, ); (3) process management (e.g., fork/exec/wait/exit(cox2011xv6, ; chickadee, ; holland2002new, ; hovemeyer2004running, ); (4) memory management (e.g., virtual memory or paging) (jos_project_list, ; CS_61_pset3, ); (5) file systems (e.g., large-file support or fine-grained synchronization) (CS_61, ; chickadee, ; JOS, ; pfaff2009pintos, ). Few labs span user–kernel boundaries or integrate multiple subsystems. In contrast, Proto’s labs are organized around target applications, so each lab naturally spans multiple subsystems and user–kernel interactions.
Student workload
In prior educational OS courses, there are typically 4–7 labs per semester, each lasting 2–3 weeks. Labs are often divided into manageable checkpoints. Checkpoints usually have dependencies and are intended to be completed in sequence. Required submission artifacts are primarily source code and design documents. Instructors typically provide starter code by removing key portions from a working implementation. The omissions range from a few essential functions (common in breadth-oriented labs (CS_61, )) to most of an entire component (common in depth-oriented labs (pfaff2009pintos, )). Proto instead emphasizes breadth through application-driven labs of wider scope: students modify tens of files across multiple components, with tasks decomposed into smaller units and arranged in non-linear dependencies (see Section 6).
2.3. Hardware Considerations
@@
We choose the Raspberry Pi 3B (Pi3) as our sole platform for several reasons. (1) Popular ISA. ARMv8 is simpler than x86, yet widely used (IoT, phones, Macs), with extensive documentation and relevance. (2) Availability. Released in 2016, Pi3 has sold over 23 million units, is priced at \$35, and will remain in production through 2029 (tomshardware2024raspberry, ). (3) Documentation. Pi3 is well-documented, with strong open-source support and an active community (raspberrypiRaspberryRaspberry, ). (4) Performance. Four Cortex-A53 cores at 1 GHz, NEON SIMD, and 1 GB of memory make Pi3 suitable for a range of gaming and media apps. (5) Peripheral Ecosystem. Pi3 is compatible with many expansion boards (called HATs (githubGitHubRaspberrypihats, )), including a Game HAT (see Section 7), which adds a 640p screen, battery, buttons, and speakers.
@@
3. Proto Overview
We next describe Proto in its complete form, which includes all the OS capabilities. This implementation is the starting point for our decomposition into a series of prototypes.
@@ 3. Proto Overview
  • donut: One or more spinning 3D tori (donut, ), comprising either textual characters (for UART text output) or pixels (for framebuffer).
  • mario: The LiteNES emulator (liteNes, ), supporting limited games including Mario Bros. (1983).
  • sysmon: A floating, transparent window that visualizes real-time CPU and memory usage.
  • slider: A slide viewer for BMP, PNG, and GIF formats, intended for the OS builders to present their design.
  • DOOM: A famous 3D game ported to virtually “anything with a screen” (itrunsdoom, ). We build on the doom generic  (doom_generic, ) project and implement key event polling and texture/renderer management. We chose not to implement sound mixing due to its complexity.
  • MusicPlayer: Intended to enhance user engagement and evoke emotions. It plays OGG files while displaying album covers. The importance of sound in games/computer programs and for tangible user experience is crucial (guillen2021role, ; kenwright2020there, ; robb2017impact, ).
  • VideoPlayer: Playback of MPEG-1 videos.
  • launcher: A GUI frontend for launching programs with an animated background.
  • blockchain: A multithreaded program for mining blocks.
Of these apps, a number of them were derived from NJU’s project-N (project-n, ). We also ported all console apps from xv6, including shell (enhanced with script execution) and utilities as ls, cat, and echo. Most user apps are in C and only a few are in C++ (blockchain).
@@
Proto is a monolithic kernel with an architecture similar to xv6. The kernel and user apps run at EL1 and EL0 of ARMv8, respectively. The OS supports user processes and both user and kernel threads, and can execute them across all four CPU cores of the Pi3. The kernel is written in a mix of C and ARMv8 assembly. It includes CPU-specific code for the ARM Cortex-A53 and board-specific device drivers for the Pi3.
Proto uses virtual memory with separate address spaces for each user app. User space starts at 0x0; kernel space uses addresses prefixed with 0xffff. The kernel maps physical memory and I/O in 1 MB blocks, while user apps use 4 KB pages. Only the user stack uses demand paging.
Proto implements 28 syscalls across three categories: task management (e.g., fork, sleep), file system (e.g., open, close), and threading/synchronization (e.g., clone, semaphore). The kernel exports device files (e.g., /dev/fb for framebuffer and /dev/events for keyboard) and proc files (e.g., /proc/cpuinfo and /proc/meminfo). We chose UNIX-like kernel interfaces so that existing apps and their dependent libraries (e.g., DOOM and SDL) can be ported with minimal modifications.
@@
Proto supports a selective set of IO peripherals, tailored for the target apps: framebuffer (for HDMI output), USB keyboards, GPIO (for Game HAT buttons), PWM (for sound output via a 3.5mm jack), SD cards, and UART. The decision to include USB is a careful balance between demonstrability and code simplicity, as discussed in Section 4.4.
Proto implements a window manager (see a photo in Figure LABEL:fig:collage(m)) allowing multiple apps to render to the screen and for input events to be dispatched to the app with focus.
@@
On the Pi3, Proto runs from an SD card with two partitions. Partition 1 contains the kernel image, which is loaded and launched by the Pi3 firmware. The kernel image packs the kernel code/data as well as an opaque ramdisk dump. The ramdisk, which is formatted with an ext2-like filesystem, incorporates all the user programs as ELF executables. Partition 2, formatted as FAT32, stores user files such as videos and game assets. At boot time, the kernel mounts its primary filesystem on the ramdisk and the secondary filesystem on the SD card’s partition 2. This setup allows users to conveniently exchange files between Proto and their personal devices (computers, phones, cameras) by copying files to and from partition 2. See Section 4.5 for details and rationale for supporting FAT32.
4. The prototypes